standfast01:

Which equation is true for the value b = 2? A. 2b + 24 = 30 B. 3b − 2 = 4 C. b + 4 = 8 D. 2b − 3 = 0

TheSmartOne:

Plug in 2 and see

TheSmartOne:

Replace b with 2 and see if you get the same thing on both sides of the equation

standfast01:

Okay

standfast01:

Okay so I still don't really get it...?

TheSmartOne:

This is what's happens when people give you answers and you think you know how to do it after that smh 2b + 24 = 30 Replace b with 2 (2*2) + 24 = 30 Simplify both sides Are the numbers on either side of the equal sign the same?

standfast01:

yeah....

TheSmartOne:

What's 2*2?? Add 24 to that Is it 30?? If you don't put any effort into your own question, you're not going to get to an answer

standfast01:

No it's 28

TheSmartOne:

Exactly, so it makes the equation false Do that for the rest of the options Plug in 2 3b - 2 = 4 (3*2) - 2 = 4 Is this true?